OpenStudy (anonymous):

Essay; show all work. Add the following polynomials: (1/10x^2 - 1/2x - 3/10) + (1/2x^2 + 1/4x - 1/5)

OpenStudy (anonymous):

1/10x^2 - 1/2x - 3/10 + 1/2x^2 + 1/4x - 1/5=

OpenStudy (anonymous):

how did you get that could you show me

jhonyy9 (jhonyy9):

1/10x2 +1/2x2 -1/2x +1/4x -3/10 -1/5 = 1+5 -2+1 -3-2 = ------ + ------ + ----- = 10x2 4x 10 = 6/10x2 -1/4x -5/10 = = 3/5x2 -1/4x -1/2 - so i think that x need to be just in numerators but indifferent the resultes are the same

jhonyy9 (jhonyy9):

right ?

OpenStudy (anonymous):

right

jhonyy9 (jhonyy9):

ok thanks good luck bye

jhonyy9 (jhonyy9):

do you understand it sure ?

OpenStudy (anonymous):

this is the answer I got if you would like to compare

OpenStudy (anonymous):

(1/10x^2 – 1/2x(-3/10))+(1/2x^2+(1/4x)(-1/5)) = (1/10)x^2 +(3/20)x + (1/2)x^2 - (1/20)x = (1/10)x^2 +(3/20)x + (5/10)x^2 - (1/20)x = (6/10)x^2 + (2/20)x =(6/10)x^2 + (1/10)x Find the product: (1/3x^2-2/3)*(1/2x -1/2) = (1/3)(x^2-2)(1/2)(x-1) = (1/6)(x^3-x^2-2x+2)